OpenStudy (anonymous):

52+2x+88+(x+10)+(x+2)=360 find x

OpenStudy (shaik0124):

140+2x+x+10+x+2=360 152+4x=360 4x=360-152 x=308/4

OpenStudy (shaik0124):

got it

OpenStudy (shaik0124):

x=77

OpenStudy (anonymous):

Yes